A ball is thrown vertically upwards with a velocity of 40m/s. Find the greatest height attained.
1
Expert's answer
2019-04-19T10:34:38-0400

We can find the maximum height attained by the ball from the kinematic equation:


"v_f^2 = v_0^2 + 2gh,"

here, "v_0 = 40 m\/s" is the initial velocity of the ball, "v_f = 0" is the final velocity of the ball at the maximum height, "g = -10 m\/s^2" is the acceleration due to gravity and "h" is the maximum height attained by the ball.

Then, we get:


"0 = v_0^2 - 2gh,""v_0^2 = 2gh,""h = \dfrac{v_0^2}{2g} = \dfrac{(40 \dfrac{m}{s})^2}{2 \cdot 10 \dfrac{m}{s^2}} = 80 m."

Answer:

"h = 80 m."